Two equal sums of money are learnt to A and B for 3 years at 4% and 6% simple interest respectively if B paid 468 rupees more than A find the sum lent each
1
Expert's answer
2019-11-26T10:09:11-0500

Principal Amount of Money


We need to find the Principal amount of money


Solution:


Amount of money lent to A = Amount of money Lent to B = P


Amount = Principal + Simple Interest =Principal + "\frac { P \times R \times T} {100}"

Amount of A =

"A_A = P + \frac { P \times R \times T} {100} = P + \frac { P \times 4 \times 3} {100}"


"A_A = P + \frac {12P} {100} = \frac {112P}{100}"

Amount of B =

"A_B = P + \frac { P \times R \times T} {100} = P + \frac { P \times 6 \times 3} {100}"

"= P + \frac {18P}{100} = \frac {118P} {100}"

B paid 468 rupees more than A


means, Amount of B = Amount of A + 468


"A_B = A_A + 468"


"\frac {118P}{100} = \frac {112P}{100} + 468"


"\frac {118P}{100} - \frac {112P}{100} = 468"

"\frac {118P - 112P} {100} = 468"


"\frac {6P} {100} = 468"

"6P = 46800"

"P = 7800"

Answer:

The sum lent each = 7800 rupees